determine the volume of oxygen needed to react competently with 2.00 L of hydrogen gas to form water vapor at constant temp & pressure
1
Expert's answer
2020-05-10T09:21:43-0400

Solution:

 2H2+O2=2H2O

"n(H_{2} )=\frac{V(H_{2} )}{V_{m} }=\frac{2}{22.4} =0.089(mol)"


0.089 mol - x mol

 2H2+O2=2H2O

2 mol - 1 mol


"x=n(O_{2})=\frac{1*0.089}{2}=0.045 (mol)\\


V(O_{2})= n(O_{2})*V_{m}=0.045*22.4=1.0(l)"

Answer:

V(O2)=1.0 l